# Benchmarking Differentially Private Tabular Data Synthesis

### v1

@@ Abstract.
Differentially private (DP) tabular data synthesis generates artificial data that preserves the statistical properties of private data while safeguarding individual privacy. The emergence of diverse algorithms in recent years has introduced challenges in practical applications, such as inconsistent data processing methods, lack of in-depth algorithm analysis, and incomplete comparisons due to overlapping development timelines. These factors create significant obstacles to selecting appropriate algorithms.
In this paper, we address these challenges by proposing a benchmark for evaluating tabular data synthesis methods. We present a unified evaluation framework that integrates data preprocessing, feature selection, and synthesis modules, facilitating fair and comprehensive comparisons. Our evaluation reveals that a significant utility-efficiency trade-off exists among current state-of-the-art methods. Some statistical methods are superior in synthesis utility, but their efficiency is not as good as most machine learning-based methods. Furthermore, we conduct an in-depth analysis of each module with experimental validation, offering theoretical insights into the strengths and limitations of different strategies.
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Private tabular data synthesis generates artificial data that preserves the statistical properties of real data while protecting individual privacy. This critical problem has a broad range of applications in practice, extending from healthcare  (Sangeetha et al., 2022; Khokhar et al., 2023) to governmental planning (Barrientos et al., 2018; Cunningham et al., 2021) and beyond. For instance, in healthcare, there is a need to share patient data for medical treatment while preserving privacy. Similarly, in government, data analysts must analyze sensitive personal attributes, such as gender and disability, while ensuring confidentiality. Addressing this challenge has thus received much attention and led to a growing research area.
@@
Differentiated privacy (DP) has become the gold standard for protecting privacy. DP ensures that the inclusion or exclusion of any single data point does not significantly affect the outcome, thereby protecting each individual data point within a dataset. A substantial body of research has been proposed to address the data synthesis problem with DP. Based on their working principles, they can be broadly classified into two categories: statistical methods and machine learning methods. Statistical methods (Zhang et al., 2017; Vietri et al., 2022; Zhang et al., 2021; Liu et al., 2023; Zhang et al., 2017; McKenna et al., 2021; Vietri et al., 2020; Hardt et al., 2012; Sun et al., 2024) compress data information through statistical properties, such as low-dimensional data distributions, to achieve data generation. On the other hand, machine learning methods leverage deep learning frameworks designed for data generation, such as generative network (Liu et al., 2021; Harder et al., 2021; Jordon et al., 2018; Xie et al., 2018; Torkzadehmahani et al., 2019; Gong et al., 2025a) and diffusion model (Kotelnikov et al., 2023; Pang et al., 2024; Li et al., 2024).
@@
In addition to these efforts to address this problem, many works also focus on providing comprehensive benchmarks. For instance, Du et al. (Du and Li, 2024) and Tao et al. (Tao et al., 2022) try to evaluate current methods under the same setting. But their works do not involve recent algorithms and thus lack completeness. Some other benchmark works (Hu et al., 2023; Yang et al., 2024; Fan, 2020) focus more on making algorithm analysis and comparison, lacking necessary empirical validation.
@@
In summary, even though several studies have been conducted in this field, we still face several challenges: (1) Lack of unified evaluation settings. Beyond algorithmic strategies, evaluation settings, such as data preprocessing, play a significant role in determining algorithm performance. However, these settings are often considered trivial and thus are frequently overlooked in many methods, which potentially leads to unfair comparisons between methods. (2) Lack of systematic and in-depth analysis. Current works often focus on proposing new methods, only providing limited intuition about algorithm analysis. Consequently, certain aspects of in-depth analysis, such as analysis for individual algorithm modules, remain underexplored. For example, questions such as how to select marginals more accurately are insufficiently addressed. However, many generative algorithms rely entirely on the selection of marginals, making this result particularly important. (3) Lack of comprehensive comparison. Due to various reasons, such as concurrent development or relatively recent introduction, comparisons between recently proposed methods and existing works remain incomplete, particularly between some representative methods like Private-GSD (Liu et al., 2023) and AIM (McKenna et al., 2022). Furthermore, current comparisons largely focus on the overall utility of algorithms, with little attention given to experiments analyzing specific working modules.
@@
In light of the above challenges, we believe proposing a new benchmark for evaluating tabular data synthesis is necessary. The contributions of our benchmark work are as follows.
@@
Providing Rigorous Analysis for Current Methods. Given our unified framework, we conduct an in-depth analysis of different modules. We consider different preprocessing methods and their drawbacks and advantages. For the feature selection module, we divide current methods into adaptive methods and non-adaptive methods based on their working principle and formally prove the superiority of introducing scale penalty term and adaptive selection strategy. Finally, we investigate the efficiencies of current synthesis methods and discuss their potential limitations.
Conducting a Comprehensive Comparison. We include current state-of-the-art methods (McKenna et al., 2022; Zhang et al., 2021; Cai et al., 2021; Vietri et al., 2022; Liu et al., 2021; Harder et al., 2021) under both statistical and machine learning methods, and newly proposed methods (Liu et al., 2023; Kotelnikov et al., 2023) that have not been thoroughly explored. Moreover, our evaluation is more fine-grained and helps us understand how each module of the algorithms functions independently.
We observed some important experimental findings, i.e., (1) A significant trade-off between utility and efficiency exists among current methods. Two statistical methods, AIM (McKenna et al., 2022) and PrivMRF (Cai et al., 2021), outperform in utility but show worse time efficiency. Machine learning-based methods, even though relatively inferior in utility, are highly efficient. (2) Preprocessing is crucial for improving algorithm efficiency without significant synthesis errors. It is also algorithm-dependent, with different techniques better suited to specific synthesis methods. (3) All current synthesis modules exhibit limitations in different ways, such as low efficiency or inferior utility. Our source code is available on GitHub. [^1] https://github.com/KaiChen9909/tab_bench
@@ 2. Problem Formulation
${\mathtt{Pr}}\left[\mathcal{A}(D)\in T\right]\leq e^{\varepsilon}\,{\mathtt{Pr}}\left[\mathcal{A}(D^{\prime})\in T\right]+\delta.$
Here, we say two datasets are neighboring ($D\simeq D^{\prime}$) when they differ on one tuple/sample. To achieve DP in different scenarios, many mechanisms have been employed, such as Gaussian mechanism (Mironov, 2017), exponential mechanism (McKenna et al., 2022; Liu et al., 2021) and DP-SGD (Mironov et al., 2019). We provide a detailed introduction to them in the appendix. In our work, we use Rényi DP as a tight composition tool, defined as:
Definition 0 (Rényi DP (Mironov, 2017)).
@@
RDP has composition and post-processing properties (Mironov, 2017), which makes it a suitable choice for complex algorithm design. Moreover, an $(\alpha,\varepsilon)$-RDP guarantee can easily be converted to a $(\varepsilon^{\prime},\delta)$-DP guarantee via Theorem 3 (Mironov, 2017).
@@
If $f$ is an $(\alpha,\varepsilon)$-RDP mechanism, then it also satisfy $\left(\varepsilon+\frac{\log{1/\delta}}{\alpha-1},\delta\right)$-DP for any $10<\delta<$ 1.
@@ 3.1. Existing Algorithms
Broadly, existing DP tabular data synthesis methods can be divided into two categories, which are statistical methods and machine learning methods.
@@
Statistical Methods. The exploration of statistical methods started earlier. Shortly after the development of DP, researchers began investigating the data generation problem under the DP. MWEM (Hardt et al., 2012) and DualQuery (Gaboardi et al., 2015) both release data by repeatedly improving an approximated distribution using Multiplicative Weight approach (Hardt and Rothblum, 2010). Another notable line of research involves Bayesian networks, such as PrivBayes (Zhang et al., 2017) and BSG (Bindschaedler et al., 2017). In addition, Li et al. (Li et al., 2014) try to utilize Copula functions for data generation.
@@
In 2018 and 2020, NIST (NIST, 2018, 2020) hosted challenges about DP data synthesis. Among the competing algorithms, PrivBayes, MST (McKenna et al., 2021), and DPSyn (Li et al., 2021) exhibited the best performance. All of these methods attempt to privately identify and answer highly correlated low-dimensional marginals. Their main differences lie in their methodology for selecting these low-dimensional marginals and in how they represent the data distribution from these noisy marginals. For example, MST synthesizes data using probabilistic graphical models (PGMs) (McKenna et al., 2019), PrivBayes uses a Bayesian model, and PrivSyn iteratively updates an initialized dataset using an algorithm they call GUM.
@@
After these NIST challenges, more advanced statistical methods were proposed. Zhang et al. proposed PrivSyn (Zhang et al., 2021) by organizing and refining DPSyn. Cai et al. introduced PrivMRF (Cai et al., 2021), and McKenna et al. proposed AIM (McKenna et al., 2022), both of which dynamically select low-dimensional marginals and employ PGMs for synthesis. Moreover, methods such as FEM (Vietri et al., 2020) and RAP/RAP++ (Aydore et al., 2021; Vietri et al., 2022) treat synthesis as an optimization problem, utilizing FTPL (Kalai and Vempala, 2005; Suggala and Netrapalli, 2020; Syrgkanis et al., 2016) and relaxed projection (Neel et al., 2018), respectively, to refine initialized datasets using adaptively selected marginals. More recently, Liu et al. (Liu et al., 2023) proposed Private-GSD, which can apply genetic algorithms to adjust datasets based on any selected marginals iteratively.
@@
Machine Learning Methods. In addition to statistical methods, machine learning models have been widely explored for DP tabular data synthesis. In NIST 2018, there was an effort to utilize GANs to generate data. However, this method, called DP-GAN (Xie et al., 2018), did not demonstrate a good performance. Further attempts on generative adversarial networks (GANs), such as DP-GAN (Xie et al., 2018), DP-WGAN (Srivastava and Alzantot, 2019), DP-CGAN (Torkzadehmahani et al., 2019), PATE-GAN (Jordon et al., 2018), and DP-CTGAN (Fang et al., 2022) also demonstrated limited performance before. Therefore, we omit the comparison of this class of methods.
@@
Statistical Method RAP (Aydore et al., 2021) Unspecified Adaptive Relaxed Projection FEM (Vietri et al., 2020),HDMM (McKenna et al., 2018)
PrivSyn (Zhang et al., 2021) Categorical Preprocessing Non-adaptive GUM PrivBayes (Zhang et al., 2017), DualQuery (Gaboardi et al., 2015),PGM (McKenna et al., 2019)
PrivMRF (Cai et al., 2021) Unspecified Adaptive PGM PrivBayes, BSG (Bindschaedler et al., 2017), DP-WGAN (Srivastava and Alzantot, 2019), DP-Copula (Li et al., 2014)
RAP++ (Vietri et al., 2022) Unspecified Adaptive Relaxed Projection RAP, DP-MERF, DP-CTGAN (Fang et al., 2022), PGM
AIM (McKenna et al., 2022) Unspecified Adaptive PGM PrivMRF, RAP, MST (McKenna et al., 2021), MWEM (Hardt et al., 2012)
Private-GSD (Liu et al., 2023) Unspecified Unspecified Genetic Algorithm GEM, RAP++, PGM (McKenna et al., 2019)
Machine Learning Method GEM (Liu et al., 2021) Unspecified Adaptive Generative Network RAP, MWEM, DualQuery
DP-MERF (Harder et al., 2021) Unspecified Non-adaptive Generative Network DP-GAN (Xie et al., 2018), DP-CGAN (Torkzadehmahani et al., 2019)
TabDDPM (Kotelnikov et al., 2023) Unspecified - Diffusion Model -

### v2

@@ Abstract.
Differentially private (DP) tabular data synthesis generates artificial data that preserves the statistical properties of private data while safeguarding individual privacy. The emergence of diverse algorithms in recent years has introduced challenges in practical applications, such as inconsistent data processing methods, the lack of in-depth algorithm analysis, and incomplete comparisons due to overlapping development timelines. These factors create significant obstacles to selecting appropriate algorithms.
In this paper, we address these challenges by proposing a benchmark for evaluating tabular data synthesis methods. We present a unified evaluation framework that integrates data preprocessing, feature selection, and synthesis modules, facilitating fair and comprehensive comparisons. Our evaluation reveals that a significant utility-efficiency trade-off exists among current state-of-the-art methods. Some statistical methods are superior in synthesis utility, but their efficiency is not as good as most deep learning-based methods. Furthermore, we conduct an in-depth analysis of each module with experimental validation, offering theoretical insights into the strengths and limitations of different strategies. Our code is open-sourced via the link. [^1] https://github.com/KaiChen9909/tab_bench
Differential Privacy, Tabular Data Synthesis
copyright: ccjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 6 (SIGMOD)article: 299publicationmonth: 12doi: 10.1145/3769764ccs: Security and privacy Privacy-preserving protocols
@@ 1. Introduction
Private tabular data synthesis generates artificial data that preserves the statistical properties of real data while protecting individual privacy. This critical problem has a broad range of applications in practice, extending from healthcare (sangeetha2022differentially; khokhar2023differentially) to governmental planning (barrientos2018providingaccessconfidentialresearch; Cunningham_2021) and beyond. For instance, in healthcare, there is a need to share patient data for medical treatment while preserving privacy. Similarly, in government, data analysts must analyze sensitive personal attributes while ensuring confidentiality. Addressing this challenge has thus received much attention.
@@
Differential privacy (DP) has become the gold standard for protecting privacy. DP ensures that the inclusion or exclusion of any single data point does not significantly affect the outcome, thereby protecting each individual data point within a dataset. A substantial body of research has been proposed to address the data synthesis problem with DP. Based on their working principles, they can be broadly classified into two categories: statistical methods and deep learning methods. Statistical methods (zhang2017privbayes; vietri2022private; zhang2021privsyn; liu2023generating; zhang2017privbayes; mckenna2021winning; vietri2020neworacleefficientalgorithmsprivate; hardt2012simple) compress data information through statistical properties, such as low-dimensional data distributions, to achieve data generation. On the other hand, deep learning methods leverage deep learning frameworks designed for data generation, such as generative adversarial networks (liu2021iterative; harder2021dp; jordon2018pate; xie2018differentially; torkzadehmahani2019dp; gong2025dpimagebench) and diffusion models (kotelnikov2023tabddpm; pang2024clavaddpm; li2024privimage).
@@
In addition to these efforts to address this problem, many works also focus on providing comprehensive benchmarks. For instance, Du et al. (du2024towards) and Tao et al. (tao2022benchmarkingdifferentiallyprivatesynthetic) try to evaluate current methods under the same setting. But their works do not involve recent algorithms and thus lack completeness. Some other benchmark works (hu2023sokprivacypreservingdatasynthesis; yang2024tabular; fan2020survey) focus more on making algorithm analysis and comparison, lacking necessary empirical validation.
@@
In summary, even though several studies have been conducted in this field, we still face several challenges: (1) Lack of unified evaluation settings. Beyond algorithmic strategies, evaluation settings, such as data preprocessing, play a significant role in determining algorithm performance. However, these settings are often considered trivial and thus are frequently overlooked in many methods, which potentially leads to unfair comparisons between methods. (2) Lack of systematic and in-depth analysis. Current works often focus on proposing new methods, only providing limited intuition about algorithm analysis, or they only provide an overall comparison without detailed analysis, such as comparing deep learning and statistical methods’ utility, but not delving into the reasons for the observed differences. Consequently, certain aspects of in-depth analysis, such as analysis for individual algorithm modules, remain underexplored. For example, questions such as how to select marginals more accurately are insufficiently addressed. However, many generative algorithms rely entirely on the selection of marginals, making this result particularly important. (3) Lack of comprehensive comparison. Due to various reasons, such as concurrent development or relatively recent introduction, comparisons between recently proposed methods and existing works remain incomplete, particularly between some representative methods like Private-GSD (liu2023generating) and AIM (mckenna2022aim). Furthermore, current comparisons largely focus on the overall utility of algorithms, with little attention given to experiments analyzing specific working modules.
@@
In light of the above challenges, we believe that proposing a new benchmark for evaluating tabular data synthesis is necessary. The contributions of our benchmark work are as follows.
@@
Providing Rigorous Analysis for Current Methods. Given our unified framework, we conduct an in-depth analysis of different modules. We consider different preprocessing methods and their drawbacks and advantages. For the feature selection module, we discuss the importance of introducing a scale penalty term during selection and prove the superiority of the adaptive marginal selection strategy. Finally, we investigate the efficiencies of current synthesis methods and discuss their potential limitations.
Conducting a Comprehensive Comparison. We include current state-of-the-art methods (mckenna2022aim; zhang2021privsyn; cai2021data; vietri2022private; liu2021iterative; harder2021dp) under both statistical and deep learning methods, and newly proposed methods (liu2023generating; kotelnikov2023tabddpm) that have not been thoroughly explored. Moreover, our evaluation is more fine-grained and helps us understand how each module of the algorithms functions independently.
We observed some important experimental findings, i.e., (1) A significant trade-off between utility and efficiency exists among current methods under current implementations. Two statistical methods, AIM (mckenna2022aim) and PrivMRF (cai2021data), outperform in utility but show worse time efficiency. Deep learning methods, even though relatively inferior in utility, are highly efficient. (2) Preprocessing is crucial for improving algorithm efficiency without significant synthesis errors. It is also algorithm-dependent, with different techniques better suited to specific synthesis methods. (3) All current synthesis modules exhibit limitations in different ways, such as low efficiency or inferior utility.
@@ 2. Problem Formulation
$\Pr{\mathcal{A}(D)\in T}\leq e^{\varepsilon}\,\Pr{\mathcal{A}(D^{\prime})\in T}+\delta.$
Here, we say two datasets are neighboring ($D\simeq D^{\prime}$) when they differ on one tuple/sample. To achieve DP in different scenarios, many mechanisms have been employed, such as Gaussian mechanism (mironov2017renyi), exponential mechanism (mckenna2022aim; liu2021iterative), and DP-SGD (mironov2019r). We provide a detailed introduction to them in the appendix of our full paper (chen2025benchmarking). In our work, we use Rényi DP as a tight composition tool, defined as:
Definition 0 (Rényi DP (mironov2017renyi)).
@@
RDP has composition and post-processing properties (mironov2017renyi), which make it a suitable choice for complex algorithm design. Moreover, an $(\alpha,\varepsilon)$-RDP guarantee can easily be converted to a $(\varepsilon^{\prime},\delta)$-DP guarantee via Theorem 3 (mironov2017renyi).
@@
If $f$ is an $(\alpha,\varepsilon)$-RDP mechanism, then it also satisfy $\left(\varepsilon+\frac{\log{1/\delta}}{\alpha-1},\delta\right)$-DP for any $0<\delta<1$.
@@ 3.1. Existing Algorithms
Broadly, existing DP tabular data synthesis methods can be divided into two categories, which are statistical methods and deep learning methods.
@@
Statistical Methods. The exploration of statistical methods started earlier. Shortly after the development of DP, researchers began investigating the data generation problem under the DP. MWEM (hardt2012simple) and DualQuery (gaboardi2015dualquerypracticalprivate) both release data by repeatedly improving an approximated distribution using Multiplicative Weight approach (hardt2010multiplicative). Another notable line of research involves Bayesian networks, such as PrivBayes (zhang2017privbayes) and BSG (bindschaedler2017plausibledeniabilityprivacypreservingdata). In addition, Li et al. (li2014differentially) try to utilize Copula functions for data generation.
@@
In 2018 and 2020, NIST (url:nistchallenge2018; url:nistchallenge2020) hosted challenges about DP data synthesis. Among the competing algorithms, PrivBayes, MST (mckenna2021winning), and DPSyn (li2021dpsynexperiencesnistdifferential) exhibited the best performance. All of these methods attempt to privately identify and answer highly correlated low-dimensional marginals. Their main differences lie in their methodology for selecting these low-dimensional marginals and in how they represent the data distribution from these noisy marginals. For example, MST synthesizes data using probabilistic graphical models (PGMs) (mckenna2019graphical), PrivBayes uses a Bayesian model, and PrivSyn iteratively updates an initialized dataset using an algorithm they call GUM.
@@
After these NIST challenges, more advanced statistical methods were proposed. Zhang et al. proposed PrivSyn (zhang2021privsyn) by organizing and refining DPSyn. Cai et al. introduced PrivMRF (cai2021data), and McKenna et al. proposed AIM (mckenna2022aim), both of which dynamically select low-dimensional marginals and employ PGMs for synthesis. Moreover, methods such as FEM (vietri2020neworacleefficientalgorithmsprivate) and RAP/RAP++ (aydore2021differentially; vietri2022private) treat synthesis as an optimization problem, utilizing FTPL (kalai2005efficient; suggala2020online; syrgkanis2016efficient) and relaxed projection (neel2018useheuristicsdifferentialprivacy), respectively, to refine initialized datasets using adaptively selected marginals. More recently, Liu et al. (liu2023generating) proposed Private-GSD, which can apply genetic algorithms to adjust datasets based on any selected marginals iteratively.
@@
Deep learning Methods. In addition to statistical methods, deep learning models have been widely explored for DP tabular data synthesis. In this work, we use the category “deep learning model” to broadly encompass all neural network-related approaches, even though some of them do not have a “deep” network structure. In NIST 2018, there was an effort to utilize GANs to generate data. However, this method, called DP-GAN (xie2018differentially), did not demonstrate good performance. Further attempts on generative adversarial networks (GANs), such as DP-GAN (xie2018differentially), DP-WGAN (srivastava2019differentially), DP-CGAN (torkzadehmahani2019dp), PATE-GAN (jordon2018pate), and DP-CTGAN (fang2022dp), also demonstrated limited performance before. Therefore, we omit the comparison of this class of methods.
@@
Statistical Method RAP (aydore2021differentially) Unspecified Adaptive Relaxed Projection FEM (vietri2020neworacleefficientalgorithmsprivate),HDMM (McKenna_2018)
PrivSyn (zhang2021privsyn) Categorical Preprocessing Non-adaptive GUM PrivBayes (zhang2017privbayes), DualQuery (gaboardi2015dualquerypracticalprivate),PGM (mckenna2019graphical)
PrivMRF (cai2021data) Unspecified Adaptive PGM PrivBayes, BSG (bindschaedler2017plausibledeniabilityprivacypreservingdata), DP-WGAN (srivastava2019differentially), DP-Copula (li2014differentially)
RAP++ (vietri2022private) Unspecified Adaptive Relaxed Projection RAP, DP-MERF, DP-CTGAN (fang2022dp), PGM
AIM (mckenna2022aim) Unspecified Adaptive PGM PrivMRF, RAP, MST (mckenna2021winning), MWEM (hardt2012simple)
Private-GSD (liu2023generating) Unspecified Unspecified Genetic Algorithm GEM, RAP++, PGM (mckenna2019graphical)
deep learning Method GEM (liu2021iterative) Unspecified Adaptive Generative Network RAP, MWEM, DualQuery
DP-MERF (harder2021dp) Unspecified Non-adaptive Generative Network DP-GAN (xie2018differentially), DP-CGAN (torkzadehmahani2019dp)
TabDDPM (kotelnikov2023tabddpm) Unspecified - Diffusion Model -
More recent deep learning approaches, including GEM (liu2021iterative) and DP-MERF (harder2021dp), represent generative network-based advancements. GEM combines generative networks with adaptive marginal selection mechanisms, while DP-MERF employs random Fourier feature loss to train generative networks. Besides, TabDDPM (kotelnikov2023tabddpm) leverages diffusion models’ representational power to fit target data directly. While TabDDPM was not originally designed for DP, it achieves state-of-the-art performance among non-DP methods and can be adapted for DP synthesis using DP-SGD (du2024towards). Consequently, we include TabDDPM in our analysis.
Our work focuses on recently proposed methods that have not been well compared previously and those representing the current state-of-the-art, as summarized in Table 1. Here, we must emphasize that specifically, we don’t consider some LLM-based methods (tran2024differentially; sablayrolles2023privatelygeneratingtabulardata). That is because LLMs are trained on extensive public datasets, which will introduce evaluation bias and affect our further comparison of algorithm modules.
3.2. Existing Benchmark Works
In addition to the proposed algorithms, there are several benchmark studies (du2024towards; hu2023sokprivacypreservingdatasynthesis; yang2024tabular; fan2020survey; tao2022benchmarkingdifferentiallyprivatesynthetic) that investigate the problem of DP data synthesis. However, previous works all have some weaknesses, summarized as follows.
  • Lack of unified comparison setting. Du et al.  (du2024towards) have made an experimental evaluation of current works, but ignore the importance of a unified setting (e.g., preprocessing), which may significantly influence the comparison fairness.
  • Not include recent advanced works. Du et al.  (du2024towards) and Tao et al. (tao2022benchmarkingdifferentiallyprivatesynthetic) focus on utility evaluation, but they both lack investigation for some advanced methods, such as RAP++ and Private-GSD. Fan et al.’s work  (fan2020survey) only focuses on GAN-based methods.
  • Lack of deep analysis. Yang et al. (yang2024tabular) provide a survey of many methods and further investigate distributed data synthesis. However, they do not delve deep into these algorithms’ working principles. Du et al.’s work and Tao et al.’s work also have a similar weakness in lacking deep algorithm analysis.
  • Lack of comprehensive experiments. Hu et al. (hu2023sokprivacypreservingdatasynthesis) provide analysis for a wide range of DP synthesis algorithms, not only about tabular data synthesis but also trajectory data and graph data. This work, though trying to make an in-depth analysis, lacks comprehensive experiments.
Except for these works, there are other “benchmark-like” works on different aspects of this problem. For instance, Ganev et al. (ganev2025importancediscretemeasuringimpact) discuss the importance of discretization in tabular data synthesis. Moreover, Stadler et al. (stadler2022syntheticdataanonymisation) focus more on the quantitative evaluation of privacy gain. These works provide in-depth research from different perspectives but lack a more comprehensive viewpoint.